# V2X-R: Cooperative LiDAR-4D Radar Fusion for 3D Object Detection with Denoising Diffusion

### v1

@@ Abstract
Current Vehicle-to-Everything (V2X) systems have significantly enhanced 3D object detection using LiDAR and camera data. However, these methods suffer from performance degradation in adverse weather conditions. The weather-robust 4D radar provides Doppler and additional geometric information, raising the possibility of addressing this challenge. To this end, we present V2X-R, the first simulated V2X dataset incorporating LiDAR, camera, and 4D radar. V2X-R contains 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes. Subsequently, we propose a novel cooperative LiDAR-4D radar fusion pipeline for 3D object detection and implement it with various fusion strategies. To achieve weather-robust detection, we additionally propose a Multi-modal Denoising Diffusion (MDD) module in our fusion pipeline. MDD utilizes weather-robust 4D radar feature as a condition to prompt the diffusion model to denoise noisy LiDAR features. Experiments show that our LiDAR-4D radar fusion pipeline demonstrates superior performance in the V2X-R dataset. Over and above this, our MDD module further improved the performance of basic fusion model by up to 5.73%/6.70% in foggy/snowy conditions with barely disrupting normal performance. The dataset and code will be publicly available at: https://github.com/ylwhxht/V2X-R.
@@ 1 Introduction
In recent years, autonomous driving and other unmanned systems have garnered widespread attention. This has led to rapid advancements in 3D object detection [24, 64, 4]. Outdoor environments, however, present complex and dynamic challenges, including various occlusions and weather conditions [44, 14]. Such factors significantly impact the performance of 3D object detection. Consequently, some effort has been made to explore multi-agent cooperative perception, such as vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-everything (V2X) [53, 52, 56, 63]. Benefiting from the information shared between agents, in complex outdoor scenarios, cooperative 3D object detection has natural advantages, such as long detection distance and multi-view object observation.
Figure 1: The advantages provided by the dense 4D radar point cloud in multi-agent view. Including weather-robustness, fewer spatial errors, Doppler information, and geometric shapes.
@@
Aside from LiDAR and camera, 4D radar has also been widely noticed [1, 16, 10, 38, 41, 40] because it can perform all-weather sensing and provide speed measurements [66]. Fusing 4D radar and LiDAR is expected to improve the perception ability under adverse weather conditions. As shown in Fig. 1, cooperative 4D radar and LiDAR fusion perception have the following advantages:
  • Weather-Robustness. The millimeter-wave signals of 4D radar easily penetrates particles in adverse weather [30, 7, 50], fusing 4D radar and LiDAR will grant the model weather robust sensing capabilities.
  • Fewer spatial relationship errors. The error-prone operations (view-transformation or depth-estimation) [43, 42, 1] are not required in the process of fusion of 4D radar and LiDAR point cloud. There are fewer corresponding spatial relationship errors in fusing 4D radar and LiDAR.
  • Additional Doppler information. The Doppler information provided by 4D radar is favorable for object detection [16, 1], which can greatly help to detect motion objects.
  • Richer geometric and shape information. The multi-agent view can significantly improve the shortcoming of low resolution in 4D radar. This empowers 4D radar with the ability to supplement the richer geometric and shape information of objects for LiDAR.
However, there is a lack of 4D radar data in the current cooperative perception dataset. Therefore, we present V2X-R, the first simulated cooperative 3D Object Detection V2X dataset that not only includes LiDAR, cameras, but also 4D radar data. V2X-R contains 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes.
Based on the V2X-R dataset, we further develop a general cooperative LiDAR-4D radar fusion pipeline for 3D object detection. The entire pipeline explicitly consists of four stages: 1) Encode by each agent. 2) Agent fusion. 3) Modal fusion. 4) Box prediction. We propose a novel Multi-modal Diffusion Denoising (MDD) module in the modal fusion stage of the pipeline to accommodate the noisy LiDAR features challenge after the agent fusion stage. MDD utilizes weather-robust 4D radar feature as a condition to prompt the diffusion model to denoise noisy LiDAR features. Notably, our MDD module barely disrupts performance in normal weather, attaining high performance in both normal and adverse weather.
@@
  • We designed a novel Multi-modal Diffusion Denoising (MDD) module to utilize weather-robust 4D radar feature to handle weather-induced noisy LiDAR features. Our MDD shows effectiveness in both single and multi-agent, simulated and real scenarios.
@@ 2.2 Cooperative Perception Datasets
Large-scale multi-sensor cooperative perception datasets, such as V2V4Real [55], are crucial for training robust and generalizable perception models. However, real-world data collection is costly and time-consuming. Simulation platforms like CARLA [6], SUMO [25], and OpenCDA [51] offer a cost-effective alternative. Datasets such as OPV2V [53] and V2X-Sim [19] have been developed using these tools, supporting tasks like 3D object detection, tracking, and bird’s-eye view (BEV) segmentation. However, the lack of 4D Radar data in existing datasets limits the research of model robustness under adverse weather conditions.
@@ 3.3 Cooperative 4D Radar Point Cloud Analysis
We analyzed the instance occupancy and points of the 4D radar point cloud before and after the multi-agent cooperative communication, as shown in Fig. 3. The 4D radar point cloud instance occupancy and points at different distances increase significantly within multi-agent collaboration. Especially in the middle and long distance, 4D radar point cloud can still have relatively high instance occupancy and points. These data demonstrate the independent perception capability of the collaborative 4D radar point cloud.
@@ 3.4 Adverse Weather Simulation
Figure 4: The pipeline of constructed cooperative LiDAR-4D radar fusion for weather-robust 3D object detection. The fusion pipeline (a) is first fed with multi-modal point cloud (LiDAR and 4D radar) from multi-agent: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Subsequent fusion is explicitly consists of four stages: 1) Encode by each agent. 2) Agent-fusion. 3)Modal-fusion. 4) Box prediction. Additionally, a Multi-modal Denoising Diffusion (b) module is integrated in the pipeline to achieve weather-robust ability. The input noisy LiDAR features are first subjected to a diffusion process, followed by $\mathcal{T}$ step denoising process with weather-robust 4D radar features as conditions to get the denoised clear LiDAR features.
@@ 4.1 Overall Statement
Built upon the V2X-R dataset, we pioneer the exploration of 3D object detection via cooperative LiDAR-4D radar fusion. To the best of our knowledge, no prior work has investigated this problem. In sec. 4.2, we introduce the first cooperative LiDAR-4D radar fusion pipeline for weather-robust 3D object detection. Furthermore, in sec. 4.3, we delve into the weather-robust cooperative LiDAR-4D radar fusion by using the weather-robustness advantage of 4D radar. We first analyze the strengths and challenges of cooperative LiDAR point cloud under adverse weather conditions. Then, we propose the Multi-modal Denoising Diffusion (MDD) module to tackle the challenges, enhancing the weather-robustness of these methods.
@@ 4.2 Fusion Pipeline
In the cooperative LiDAR-4D radar fusion pipeline, there are three types of agents: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Each agent collects LiDAR and 4D radar point cloud data, forming the multi-agent multi-modal input $\mathbf{X}=\{X^{L}_{C},X^{L}_{E},X^{L}_{I},X^{R}_{C},X^{R}_{E},X^{R}_{I}\}$. As illustrated in the Fig. 4(a), subsequent fusion is explicitly consists of four stages:
@@
(a) Single-Agent Multi-Modal to Multi-Agent Multi-Modal (SA2MA). LiDAR-4D radar fusion in single-agent has received considerable attention [41, 40, 1, 16]. Therefore, we extend existing single-agent LiDAR-4D radar fusion methods to the multi-agent version by integrating a self-attention based agent-fusion module in the 2nd stage.
(b) Multi-Agent Single-Modal to Multi-Agent Multi-Modal (SM2MM). Another approach is to extend the widely researched multi-agent LiDAR-based methods [12, 31, 53, 54, 52] to the multi-modal version. Specifically, we first extract multi-agent features from LiDAR and 4D radar point clouds individually and then concatenate BEV features in multi-modal fusion of the 3rd stage.
@@ 4.3 Multi-modal Denoising Diffusion (MDD)
$\delta_{gau}\sim\mathcal{N}(\sqrt{\overline{\alpha}_{\mathcal{T}}}\delta_{raw},\sqrt{1-\overline{\alpha}_{\mathcal{T}}})$ subject to Gaussian distribution, which facilitates the fitting of the denoising model $U_{\theta}$. Subsequently, at each denoising step, we incorporate weather-robust 4D radar features $\mathcal{F}^{R}_{\mathcal{A}}$ as a condition to prompt U-Net [34] denoiser $U_{\theta}$ to predict the denoised feature:
@@
For the final output after $\mathcal{T}$ denoising steps, the $U_{\theta}$ generates the denoised clear LiDAR feature $\tilde{\mathcal{F}^{L}_{\mathcal{A}}}$. We present the detailed training and inference pipelines MMD in Algorithm 1. And we compute the loss of MDD as :
@@ 4.4 Loss Function
where $\beta_{cls}$, $\beta_{loc}$ are hyper-parameters, the $\mathcal{L}_{cls}$ and $\mathcal{L}_{loc}$ represents classification and localization loss, respectively. All hyper-parameters are detailed in the supplementary material.
@@ 5.2 Benchmark Models
Table 3: Results of various cooperative LiDAR-based methods on the validation and testing of our V2X-R dataset in different IoU (0.3,0.5,0.7). The representations of $\dagger$ and $\ddagger$ as described in 5.2.
@@
Table 5: The 3D mAP performance comparison under different weather conditions on the V2X-R datatset. The best performance for each model is shown as bold. ’L’ and ’4DR’ represent LiDAR and 4D radar, respectively. ’-’ indicates that multi-modal method L4DR [16] cannot achieve LiDAR-based performance.
@@
To intuitively compare the performance in normal weather differences between modalities, the cooperative LiDAR-4D radar fusion 3D object detectors we implemented in the benchmark section do not incorporate the MDD module. The MDD module will be discussed in subsequent experiments. Furthermore, other detailed implementation configurations of the selected benchmark model are described in the supplementary material.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
PointPillars 69.04 70.72 66.75
Sedan PointPillars w/ MDD 74.03 75.90 67.18
PointPillars 50.95 48.78 53.33
Bus PointPillars w/ MDD 54.58 54.75 52.81
Table 6: 3D mAP performance comparison on K-Radar [28] dataset. Adverse represents average results under various adverse weather, including overcast, fog, rain, sleet, light and heavy snow.
Performance comparison under different real-world weather on K-Radar dataset. Although simulated adverse weather performance is consistent with real-world results to some extent, a gap still exists between them. We conducted experiments on the K-Radar single-agent real-world dataset using PointPillars, the most common architecture in multi-agent cooperative perception, as the baseline. As shown in Table 6, our MDD module demonstrates significant performance improvements in real-world adverse weather, with 3D mAP gains of 5.20% and 5.97% for the Sedan and Bus classes, respectively. Consistent with the V2X-R results, the MDD module has rarely impacted the normal weather performance, resulting in overall improvements of 4.99% and 3.63%. These results validate the effectiveness of the MDD module in the presence of dense noise under adverse weather, whether it is single or multi-agent perception.
@@
Effect of diffusion layers and loss weight. Finally, we investigate the impact of diffusion layers $\mathcal{T}$ and the loss weight of MDD in our MDD module. Table 8 presents representative results. We first explore the MDD loss weight with $\mathcal{T}=2$, considering both constant weights and the epoch-varying weight $\gamma(e,\psi)$ in eqs. 9. The best performance is achieved with $\gamma(e,\psi=3)$ when $\mathcal{T}=2$. We further experiment with different values of $\mathcal{T}$. When $\mathcal{T}=2$ and the loss weight is $\gamma(e,\psi=3)$, our MDD module achieves the best performance of 87.37/83.90/86.64 for 3DmAP@Fog. Although MDD inevitably introduces an additional inference time of 32 ms, it significantly improves weather robustness and still maintains real-time (about 20 FPS). These results validate the effectiveness of our designed weight $\gamma(e,\psi)$ and real-time of our MDD module.
@@ 6 Conclusion
We present V2X-R, the first simulated V2X collaborative perception dataset with 4D radar. We establish a benchmark on V2X-R dataset by integrating the LiDAR-4D radar fusion pipeline into classic collaborative 3D object detection methods. Moreover, we propose the MDD module to address dense noise in collaborative conditions. In summary, V2X-R dataset and our pioneer exploration pave the way for collaborative LiDAR and 4D radar fusion perception.

### v2

# V2X-R: Cooperative LiDAR-4D Radar Fusion with Denoising Diffusion for 3D Object Detection

@@ Abstract
Current Vehicle-to-Everything (V2X) systems have significantly enhanced 3D object detection using LiDAR and camera data. However, they face performance degradation in adverse weather. Weather-robust 4D radar, with Doppler velocity and additional geometric information, offers a promising solution to this challenge. To this end, we present V2X-R, the first simulated V2X dataset incorporating LiDAR, camera, and 4D radar modalities. V2X-R contains 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes. Subsequently, we propose a novel cooperative LiDAR-4D radar fusion pipeline for 3D object detection and implement it with multiple fusion strategies. To achieve weather-robust detection, we additionally propose a Multi-modal Denoising Diffusion (MDD) module in our fusion pipeline. MDD utilizes weather-robust 4D radar feature as a condition to guide the diffusion model in denoising noisy LiDAR features. Experiments show that our LiDAR-4D radar fusion pipeline demonstrates superior performance in the V2X-R dataset. Over and above this, our MDD module further improved the foggy/snowy performance of the basic fusion model by up to 5.73%/6.70% and barely disrupting normal performance. The dataset and code will be publicly available at: https://github.com/ylwhxht/V2X-R.
@@ 1 Introduction
Autonomous driving and other unmanned systems have garnered widespread attention in recent years. This has led to rapid advancements in 3D object detection [24, 65, 4]. Outdoor environments, however, present complex and dynamic challenges, including various occlusions and weather conditions [45, 14]. Such factors significantly impact the performance of 3D object detection. Consequently, some effort has been made to explore multi-agent cooperative perception, such as vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-everything (V2X) [54, 53, 57, 64]. Benefiting from the information shared between agents, in complex outdoor scenarios, cooperative 3D object detection has natural advantages, such as long detection distance and multi-view object observation.
Figure 1: The advantages of the dense 4D radar point cloud in multi-agent view. Including weather robustness, fewer spatial errors, Doppler information, and geometric shapes.
@@
Aside from LiDAR and camera, 4D radar has also been widely noticed [1, 16, 10, 39, 42, 41] because it can perform all-weather sensing and provide speed measurements [67]. Fusing 4D radar and LiDAR is expected to improve the perception performance under adverse weather conditions. As shown in Fig. 1, cooperative 4D radar and LiDAR fusion perception have the following advantages:
  • Weather robustness. The millimeter-wave signals of 4D radar easily penetrates particles in adverse weather [31, 7, 51]. Fusing 4D radar and LiDAR will grant the model weather robust sensing capabilities.
  • Fewer spatial relationship errors. The error-prone operations (view-transformation or depth-estimation) [44, 43, 1] are not involved in the process of 4D radar and LiDAR point cloud fusing. Thus, there are fewer corresponding spatial relationship errors in fusing 4D radar and LiDAR.
  • Additional Doppler information. The Doppler information provided by 4D radar is favorable for object detection [16, 1], significantly aiding in the detection of moving objects.
  • Richer geometric shape information. The multi-agent view can significantly address the limitation of low resolution in 4D radar. This empowers 4D radar’s capacity to complement LiDAR with more geometric shape information of objects.
However, there is a lack of 4D radar data in the current cooperative perception dataset. Therefore, we present V2X-R, the first simulated cooperative 3D Object Detection V2X dataset that not only includes LiDAR, cameras, but also 4D radar data. V2X-R contains 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes. Built upon this dataset, we develop a general cooperative LiDAR-4D radar fusion pipeline for 3D object detection. The entire pipeline consists of four stages: 1) Encode by each agent. 2) Agent fusion. 3) Modal fusion. 4) Box prediction. To address the challenge of agent-fused LiDAR features becoming noisy in adverse weather, we propose a novel Multi-modal Diffusion Denoising (MDD) module in the modal fusion stage of the pipeline. MDD transforms the noise feature distribution into the easy-to-fit Gaussian distribution by reparameterization, which solves the challenge of complex and variable weather noise features that are difficult to fit. It further utilizes weather-robust 4D radar feature as a condition to prompt the diffusion model to denoise noisy LiDAR features. Notably, our MDD module barely disrupts performance in normal weather, attaining high performance in both normal and adverse weather.
@@
  • We designed a novel Multi-modal Diffusion Denoising (MDD) module to utilize reparameterization and weather-robust 4D radar feature to handle hard-to-fit noisy LiDAR feature. Our MDD shows effectiveness in both single and multi-agent, simulated and real scenarios.
@@ 2.2 Cooperative Perception Datasets
Large-scale multi-sensor cooperative perception datasets, such as V2V4Real [56], are crucial for training robust and generalizable perception models. However, real-world data collection is costly and time-consuming. Simulation platforms like CARLA [6], SUMO [25], and OpenCDA [52] offer a cost-effective alternative. Datasets such as OPV2V [54] and V2X-Sim [19] have been developed using these tools, supporting tasks like 3D object detection, tracking, and bird’s-eye view (BEV) segmentation. However, above datasets lack of 4D Radar data, which limits the research of model robustness under adverse weather conditions.
@@ 3.3 Cooperative 4D Radar Point Cloud Analysis
We analyzed the instance occupancy and points of the 4D radar point cloud before and after the multi-agent cooperative communication, as shown in Fig. 3. The 4D radar point cloud instance occupancy and points at different distances increase significantly with multi-agent collaboration. Especially in the middle and long distance, 4D radar point clouds can still have relatively high instance occupancy and points. These data demonstrate the independent perception capability of the cooperative 4D radar point cloud.
@@ 3.4 Adverse Weather Simulation
Figure 4: The pipeline of constructed cooperative LiDAR-4D radar fusion for weather-robust 3D object detection. The fusion pipeline (a) is first fed with multi-modal point cloud (LiDAR and 4D radar) from multi-agent: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Subsequent fusion consists of four stages: 1) Encode by each agent. 2) Agent fusion. 3)Modal fusion. 4) Box prediction. Additionally, a Multi-modal Denoising Diffusion (b) module is integrated into the pipeline to achieve weather-robust ability. The input noisy LiDAR features are first subjected to a diffusion process, followed by $\mathcal{T}$ step denoising process with weather-robust 4D radar features as conditions to get the denoised clear LiDAR features.
@@ 4.1 Overall Statement
Built upon the V2X-R dataset, we present the first exploration of cooperative LiDAR-4D radar fusion for 3D object detection. To the best of our knowledge, no prior work has investigated this problem. In sec. 4.2, we introduce the first cooperative LiDAR-4D radar fusion pipeline for weather-robust 3D object detection. Furthermore, in sec. 4.3, we enhance the weather-robustness of cooperative LiDAR-4D radar fusion by leveraging the advantage of 4D radar in adverse weather conditions. We analyze the strengths and challenges of cooperative LiDAR point clouds in adverse weather conditions and propose the Multi-modal Denoising Diffusion (MDD) module to tackle these challenges, significantly improving the pipeline’s weather-robustness.
@@ 4.2 Fusion Pipeline
In the cooperative LiDAR-4D radar fusion pipeline, there are three types of agents: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Each agent collects LiDAR and 4D radar point cloud data, forming the multi-agent multi-modal input $\mathbf{X}=\{X^{L}_{C},X^{L}_{E},X^{L}_{I},X^{R}_{C},X^{R}_{E},X^{R}_{I}\}$. As illustrated in the Fig. 4(a), subsequent fusion consists of four stages:
@@
(a) Single-Agent Multi-Modal to Multi-Agent Multi-Modal (SA2MA). LiDAR-4D radar fusion in single-agent has received considerable attention [42, 41, 1, 16]. Therefore, we extend existing single-agent LiDAR-4D radar fusion methods to the multi-agent version by integrating a self-attention-based agent-fusion module (2nd stage).
(b) Multi-Agent Single-Modal to Multi-Agent Multi-Modal (SM2MM). Another approach is to extend the widely researched multi-agent LiDAR-based methods [12, 32, 54, 55, 53] to the multi-modal version. Specifically, we first extract multi-agent features from LiDAR and 4D radar point clouds individually and then concatenate BEV features in multi-modal fusion (3rd stage).
@@ 4.3 Multi-modal Denoising Diffusion (MDD)
$\delta_{gau}\sim\mathcal{N}(\sqrt{\overline{\alpha}_{\mathcal{T}}}\delta_{raw},\sqrt{1-\overline{\alpha}_{\mathcal{T}}})$ subject to Gaussian distribution, which facilitates the fitting of the denoising model. Subsequently, at each denoising step, we incorporate weather-robust 4D radar features $\mathcal{F}^{R}_{\mathcal{A}}$ as a condition to prompt U-Net [35] denoiser $U_{\theta}$ to predict the denoised feature:
@@
For the final output after $\mathcal{T}$ denoising steps, the $U_{\theta}$ generates the denoised clear LiDAR feature $\tilde{\mathcal{F}^{L}_{\mathcal{A}}}=\mathcal{F}_{0}$. We present the detailed training and inference pipelines MMD in Algorithm 1. And we compute the loss of MDD as :
@@ 4.4 Loss Function
where $\beta_{cls}$, $\beta_{loc}$ are hyper-parameters, $\mathcal{L}_{cls}$ and $\mathcal{L}_{loc}$ are classification and localization loss, respectively. All hyper-parameters are detailed in the supplementary material.
@@ 5.2 Benchmark Models
Table 3: Experimental 3D object detection results of various cooperative LiDAR-based methods on the validation and testing of our V2X-R dataset in different IoU (0.3,0.5,0.7). The representations of $\dagger$ and $\ddagger$ as described in 5.2.
@@
Table 5: The 3D mAP performance comparison under different weather conditions on the V2X-R dataset. ’L’ and ’4DR’ represent LiDAR and 4D radar, respectively. ’-’ indicates that multi-modal method L4DR [16] cannot achieve LiDAR-based performance.
@@
To intuitively compare the performance in normal weather differences between modalities, the cooperative LiDAR-4D radar fusion 3D object detectors we implemented in the benchmark section do not incorporate the MDD module. The MDD module will be discussed in subsequent experiments.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
AttFuse [54] 69.04 70.72 66.75
Sedan AttFuse w/ MDD 74.03 75.90 67.18
AttFuse [54] 50.95 48.78 53.33
Bus AttFuse w/ MDD 54.58 54.75 52.81
Table 6: The 3D mAP performance comparison on K-Radar [28] dataset. Adverse represents average results under various adverse weather, including overcast, fog, rain, sleet, light and heavy snow. represents AttFuse [54] is implemented as single agent method by removing the agent fusion module.
Performance comparison under different real-world weather on K-Radar dataset. We further conducted experiments on the K-Radar single-agent real-world dataset. As shown in Table 6, our MDD module demonstrates significant performance improvements in real-world adverse weather, with 3D mAP gains of 5.20% and 5.97% for the Sedan and Bus classes, respectively. Consistent with the results on V2X-R dataset, MDD has rarely impacted the normal weather performance, resulting in overall improvements of 4.99% and 3.63%. These results validate the effectiveness of the MDD module in the presence of dense noise under adverse weather, whether single or multi-agent perception. Detailed evaluation metrics and performance of each weather can be found in the supplementary materials.
@@
Effect of diffusion layers and loss weight. Finally, we investigate the impact of diffusion layers $\mathcal{T}$ and the loss weight of MDD in our MDD module. Table 8 presents representative results. We first explore the MDD loss weight with $\mathcal{T}=2$, considering both constant weights and the epoch-varying weight $\gamma(e,\psi)$ in eqs. 9. The best performance is achieved with $\gamma(e,\psi=3)$ when $\mathcal{T}=2$. We further experiment with different values of $\mathcal{T}$. When $\mathcal{T}=2$ and the loss weight is $\gamma(e,\psi=3)$, our MDD module achieves the best performance of 87.37/83.90/86.64 3DmAP@Fog. Although MDD inevitably introduces an additional inference time of 32 ms, it significantly improves weather robustness and still maintains real-time (about 20 FPS). These results validate the effectiveness of our designed weight $\gamma(e,\psi)$ and real-time of our MDD module.
@@ 6 Conclusion and Discussion
We present V2X-R, the first simulated V2X collaborative perception dataset incorporating 4D radar. We establish a benchmark on V2X-R dataset by integrating a LiDAR-4D radar fusion pipeline into classic collaborative 3D object detection methods. Moreover, we propose the MDD module to tackle dense noise in collaborative conditions. In summary, V2X-R and our exploration paves the way for advancing collaborative LiDAR and 4D radar fusion perception.